# Instance-level Performance Prediction for Long-form Generation Tasks

### v1

@@ 5.2 Evaluating Performance Prediction
$f_{t,k}$ (Sec. 3) must return both a point estimate $\hat{s}$ and a predictive distribution (from which we take a $(1-\alpha)$ interval), and we evaluate these separately.
Point estimates. We report RMSE between prediction $\hat{s}$ and gold score $q$, providing an interpretable measure of average error magnitude in the target’s original units and serving as our primary evaluation metric. Pearson Correlation (PC) is shown only as a diagnostic for feature usefulness (Appendix 9); it does not directly reflect absolute error, and correlation may be high while still incurring large systematic errors. However, as it is unscaled, it is useful for comparing between tasks and metrics.
Prediction intervals. Expected Calibration Error (ECE) is popular for classification, measuring how well predicted probabilities match observed frequencies via binning (Naeini et al., 2015). However, ECE is designed for binary scores, not for continuous ones. Prediction intervals for continuous scores can be visualized in Figure 1, showing each instance’s true score on the y-axis as well as the prediction regressor on the x-axis. Ideally, predictions are not too narrow (overconfident) or too wide (underconfident), and several metrics exist for assessing such miscalibrations.
@@
We use the Continuous Ranked Probability Score (CRPS) (Gneiting and Raftery, 2007) as our primary metric for evaluating prediction intervals. It is a proper scoring rule that jointly rewards calibration and sharpness of the full predictive CDF $F$, where lower CRPS is better:
@@
For interpretability, we also measure Average Coverage Error (ACE) (Shrivastava et al., 2016; Xie et al., 2023), which compares empirical coverage of the nominal $(1-\alpha)$ interval to its target as $\mathrm{ACE}=\bigl{|}\mathrm{PICP}(p)-p\bigr{|}$ where
@@
and $p=1-\alpha$ (we use $p=0.95$). Lower ACE indicates better calibration, $p>0$ indicates the prediction is under-confident, and $p<0$ indicates it is over-confident. Note that one disadvantage of ACE is that it does not penalize excessively wide intervals (e.g., under-confident intervals that always cover $s_{i}$ incur a maximum $0.05$ penalty when $p=0.95$) as much as narrow ones. Hence, we treat ACE as a secondary, human-readable diagnostic and rely on CRPS for model selection.
6 Results
In this section, we address three central questions: 1) Is instance-level performance prediction of continuous, reference-based metrics for long-form LLM outputs feasible? 2) How much supervision is necessary to achieve reasonable performance? and 3) How do confidence-based regression and reference-free LLM-as-a-Judge (Section 4) compare? We evaluate all baselines across 11 datasets and multiple metrics (Section 5.1) using RMSE to judge point estimate accuracy and CRPS to judge calibration of predictive distributions, with ACE and PC used for additional diagnostics (Section 5.2). Our key findings are as follows.
Performance prediction is indeed feasible: regression on off-the-shelf confidence estimates achieves an average RMSE of 0.14 and CRPS of 0.08, with intervals calibrated to within 16% of nominal coverage and requiring as few as 16 training examples to converge. Tables 1 and 2 show that regression on the consistency-based “DegMat” confidence estimate performs best in terms of both RMSE and CRPS.
@@ 6 Results
Only 16 training instances are required for the CE-Reg’s test loss to stabilize. As long-form gold references can be expensive to collect, it is important to understand how sample-efficient these performance prediction methods would be. To answer that, we vary the size of the training set from as few as 4 instances and doubling up to the full training data size for CE-Reg. For RF-LLMaaJ we limit up to 16 instances. CE-Reg achieves close to our best observed accuracy with remarkably few labeled examples: Figure 2 shows that RMSE converges after just 16 training instances across all datasets. Even when matched to RF-LLMaaJ’s maximum of 16 in-context examples (Table 4), CE-Reg delivers substantially lower error. Although performance varies slightly across datasets, the convergence point is consistent, demonstrating that this regression approach to performance prediction can be effectively trained with a modest sample size, making it easy to collect training data even for out-of-domain tasks and mitigating potential domain differences.
CE-Reg outperforms RF-LLMaaJ. In Tables 1 and 2, CE-Reg methods achieve RMSEs averaging in the 0.14–0.15 range and reaching 0.35 at worst, depending on the task metric. Meanwhile, RF-LLMaaJ incurs much larger errors (averaging around 0.29, 0.85 at worst). Similarly, CE-Reg attains CRPS values between 0.07–0.12 averaged over task metrics, compared to about 0.21 for RF-LLMaaJ. Interestingly, most methods exhibit an over-confidence (intervals are too narrow) in ACE (95%) (Appendix 8), with verbalized CE-Reg producing ACEs of –0.89 and –0.79 and RF-LLMaaJ yielding ACEs of –0.51 and –0.52, while only some of the consistency-based features come close to nominal coverage (e.g. DegMat ACE: –0.16). Notably, this pattern largely reflects our use of RandomForest regression: substituting a XGBoost model markedly improves ACE, albeit at the expense of other performance metrics (see Figure 3 in Appendix). In addition, Judge scores and human ratings in the BiGGen-Bench-Results dataset yield similar performance, indicating our baselines generalize beyond evaluation metrics.
Prediction error is primarily metric-dependent. Across most tasks and methods, CodeBLEU and BERTScore are the easiest metrics to predict: CE-Reg consistently achieves low RMSE and CRPS and high PC on these metrics, and RF-LLMaaJ exhibits a similar trend on most datasets. The G-Eval in text summarization tasks are consistently low in RMSE and CRPS but also low in PC – this can be explained by the fact that its target score distribution is tightly concentrated. CodeJudge and LLMEvals tend to be the most difficult with high RMSE and CRPS and relatively low PC. Variation in RMSE and CRPS across evaluation metrics shows that prediction difficulty is largely metric dependent, with low metric-to-metric variability within each task. In our findings, we do not observe a similar trend for task-to-task variability. Therefore, we do not conclude anything about relative inherent difficulty between tasks.
@@
Graph-based consistency features consistently outperform all other CE-Reg methods across every dataset and metric Among all CE-Reg, the four graph-based consistency measures (DegMat, Eccentricity, EigVal, LexSim) achieve the lowest CRPS, indicating that capturing variability across multiple generations provides a stronger predictive signal than one- or two-step verbalized confidence prompts, on all types of dataset or metrics. By contrast, the NumSemSets consistency metric performs on par with the verbalized methods, trailing the graph-based techniques. Better results from graph-based consistency over NumSemSets and verbalized confidence are more obviously reflected in the ACE and PC measures (Tables 8 and 9).
RF-LLMaaJ is a poor performance predictor. Both point prediction and interval estimation are ill-calibrated compared to CE-Reg methods. In fact, only on a few tasks does RF-LLMaaJ seem to have any ability to predict evaluation metrics at all – Appendix Table 9 shows near-zero correlation between predicted and actual scores for the majority of task metrics. Furthermore, increasing in-context examples from 4 to 16 has negligible benefit. There are some task metrics where RF-LLMaaJ significantly better than random chance, but the variability across tasks and metrics highlights how sensitive and unreliable this approach can be, relative to the consistent performance seen in the consistency-based confidence regression approach. One reason that RF-LLMaaJ struggles to predict evaluation metrics could be its sensitivity to the prompt. A key insight is that, despite both using in-prompt LLM predictions, verbalized confidence is overall more reliable than RF-LLMaaJ, which is prompted to predict arbitrary task metrics. This may reflect the LLM’s pre-training data, which likely contains the general notion of confidence much more so than task-specific metrics like CodeBLEU or GEMBA; and the few-shot examples are not enough to overcome this deficiency.
@@
Table 3: Direct Assessment: For BiGGen we evaluate two baselines using pre-generated outputs—Verbalized 2S and RF-LLMaaJ—for directly predicting human ratings. Consistency-based (requiring additional outputs) and the Verbalized 1S (requiring prompt re-design) are omitted.
@@
RMSE 0.17 0.18 0.17 0.23 0.22 0.22
CRPS 0.10 0.10 0.10 0.14 0.14 0.14
Table 4: Comparison of RMSE and CRPS for CE-Reg using DegMat feature versus reference-free LLM-as-a-Judge (RF-LLMaaJ) using different number of training examples.
(a) RMSE vs. Training Data Instances
(b) CRPS vs. Training Data Instances
Figure 2: Impact of Training Data Size on CE-Reg Performance and Test Set RMSE and CRPS Across Datasets, using the selected task-specific metric (see Table 6). The plot shows RMSE and CRPS values for training data sizes capped at 64 instances. Notably, both stabilize after 16 instances, suggesting a performance plateau for certain datasets
@@ 7 Conclusion
This work introduces the task of instance-level performance prediction for long-form LLM generation, in which a black-box model’s input–output pair is used to predict arbitrary continuous, reference-based metric scores with associated prediction intervals. To support progress, we release a comprehensive benchmark spanning 11 datasets across summarization, machine translation, code generation, and question-answering, with multiple task evaluation metrics to predict.
Our experiments compare two families of performance prediction approaches as baselines. The first–confidence-based regression–extracts off-the-shelf verbalized and consistency-based confidence signals, then trains classic regression models to output both point estimates and prediction intervals. The second–a reference-free LLM-as-a-Judge–asks an LLM, via in-context prompting, to yield metric estimates and 95% intervals for each input-output-instance. We find that confidence-based regression, particularly when using consistency features, yields substantially better RMSE and CRPS than our implemented LLM-based judging. This demonstrates that lightweight supervised regression of simple confidence estimates can outperform few-shot LLM judging for both accuracy and uncertainty quantification.
Our work lays the groundwork for research in uncertainty-aware, per-instance continuous-scored performance prediction. We hope our benchmark, code, and baselines will encourage community efforts toward robust, uncertainty-aware evaluation and deployment of long-form LLMs.
@@ Acknowledgments
This research was supported in part by Cisco, Amazon, and by Good Systems [^1] https://goodsystems.utexas.edu/ (a UT Austin Grand Challenge dedicated to developing responsible AI technologies). We thank the Texas Advanced Computing Center (TACC) for use of its superb infrastructure. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
@@ A.2 Sample Prompts
      You will be given a news article. You will then be given one summary written for this article.        Your task is to rate the summary on one metric.        Please make sure you read and understand these instructions carefully.    Please keep this document open while reviewing, and refer to it as needed.        Evaluation Criteria:        {evaluation_instructions}        Evaluation Steps:        1. Read the news article carefully and identify the main facts and key points.    2. Read the summary and compare it to the article. Check if the summary meets the {metric} criterion.    3. Assign a score for {metric} based on the Evaluation Criteria.        Source Text:        {source}        Summary:        {prediction}            Evaluation Form (scores ONLY), provide your response in the following format:    - {metric}:        evaluation_criteria = {        "Coherence": """Coherence (1-5) - the collective quality of all sentences. A well-structured summary should build from sentence to a coherent body of information about a topic.""",        "Consistency": """Consistency (1-5) - the factual alignment between the summary and the        source document. A factually consistent summary contains only statements entailed by the source document and avoids hallucinated facts.""",        "Fluency": """Fluency (1-3): the quality of the summary in terms of grammar, spelling, punctuation, word choice, and sentence structure.            - 1: Poor. The summary has many errors that make it hard to understand or sound unnatural.        - 2: Fair. The summary has some errors that affect the clarity or smoothness of the text, but the main points are still comprehensible.        - 3: Good. The summary has few or no errors and is easy to read and follow.""",        "Relevance": """Relevance (1-5) - selection of important content from the source. The summary should include only important information from the source document. Annotators were instructed to penalize summaries which contained redundancies and excess information."""    }
@@
      Task: Score the following LLM output of a question-answering task with respect to the following aspects using a 1 to 5 star rating system.            Dataset: The dataset is a Question-Answering dataset, specifically designed for evaluating factual precision and detailed comparative reasoning in AI-generated answers.            Output: Begin your evaluation by providing a short explanation. Be as objective as possible. After your explanation, provide your scores in JSON format like: [[SCORE]] {{"accuracy": 2, "informativeness": 3}}            Criteria:        - Accuracy:            1 star: Incorrect information            2 stars: Partially correct information            3 stars: Half correct information            4 stars: Mostly correct information            5 stars: Perfectly correct information            - Informativeness:            1 star: No or irrelevant information            2 stars: Very little information            3 stars: Some information            4 stars: Enough information            5 stars: Highly informative            Question:        {source}            Provided Answer:        {ans}            Reference Answer(s):        {ref_text}            Evaluation:        [[SCORE]] {{\"accuracy\": <1–5>, \"informativeness\": <1–5>}}
@@
      You are a fair judge assistant tasked with providing clear, objective feedback based on specific criteria, ensuring each assessment reflects the absolute standards set for performance.
@@
      ###Task Description:    An instruction (might include an Input inside it), a response to evaluate, and a score rubric representing an evaluation criteria are given.    1. Write a detailed feedback that assesses the quality of the response strictly based on the given score rubric, not evaluating in general.    2. After writing a feedback, write a score that is an integer between 1 and 5. You should refer to the score rubric.    3. The output format should look as follows: "(write a feedback for criteria) [RESULT] (an integer number between 1 and 5)"    4. Please do not generate any other opening, closing, and explanations.        ###The instruction to evaluate:    {instruction}        ###Response to evaluate:    {response}        ###Score Rubrics:    {rubric}        ###Feedback:
@@
      You will be provided with a problem statement, a code snippet that supposedly addresses the problem,    and a catalog of code inconsistencies.    Evaluation Steps:    1. Read the problem statement carefully to identify the functionalities required for the    implementation.    2. Read the code snippet and compare it to the problem statement. Check if the code snippet covers    the required functionalities.    3. Output your answer in a JSON format list.    a) If the code snippet is correct, output: ["inconsistency": "None", "severity": "Negligible"].    b) If the code snippet is incorrect, output the identified inconsistencies and their severity    according to the catalog of code inconsistencies. For example: ["inconsistency": "<inconsistency1>",    "severity": "<severity1>", "inconsistency": "<inconsistency2>", "severity": "<severity2>", ...]    Problem: {PROBLEM}    Code Snippet: {CODE}    Taxonomy of Common Inconsistencies:    1. Missing dependency declarations: Negligible    2. No error messages for unexpected input cases: Negligible    3. Inefficiency, unnecessary statements: Negligible    4. Edge case not handled: Small    5. Logic error: Major    6. Function or variable not defined: Fatal    7. Code not completed: Fatal    Evaluation Form:    JSON output (a JSON list only):

### v2

@@ 5.2 PIE Evaluation
Predictor $f_{t,k}$ (Section 3) must return both a point estimate $\hat{s}$ and a predictive distribution (with a $(1-\alpha)$ interval). We evaluate these with 4 PIE metrics (not to be confused with task-specific metrics).
@@
Continuous Ranked Probability Score (CRPS) (Gneiting and Raftery, 2007) is PIE’s primary metric for evaluating both point estimates and prediction intervals. It is a proper scoring rule that jointly rewards calibration and sharpness of the full predictive CDF $F$; lower CRPS is better:
@@
Prediction intervals for continuous scores can be visualized in Figure 1, showing each instance’s true score on the y-axis and the prediction regressor on the x-axis. Ideally, predictions are neither too narrow (overconfident) nor too wide (underconfident). While Expected Calibration Error (ECE) is widely used for classification to compare predicted probabilities with observed frequencies via binning (Naeini et al., 2015), it is traditionally defined over binary scores, not continuous ones. Kuleshov et al. (2018) propose an ECE variation for continuous scores based on quantile coverage, but unlike CRPS, such diagnostics do not reward distribution sharpness and can therefore favor overly conservative (wide) intervals (Gneiting et al., 2007).
Other complementary metrics We also report RMSE (Appendix 8), providing an interpretable measure of average error magnitude in the target’s original units. Pearson Correlation (PC) is also used as a diagnostic for feature usefulness (Appendix 9); it does not directly reflect absolute error, and correlation may be high while still incurring large systematic errors. However, as it is unscaled, it is useful for comparing between tasks and metrics and for reflecting instance-level discriminative predictability.
In addition, we also measure Average Coverage Error (ACE) (Shrivastava et al., 2016; Xie et al., 2023) (Appendix 10), which compares empirical coverage of the nominal $(1-\alpha)$ interval to its target as $\mathrm{ACE}=\bigl|\mathrm{PICP}(p)-p\bigr|$ where
@@
and $p=1-\alpha$ (we use $p=0.95$). Lower ACE indicates better calibration, with positive $p$ reflecting under-confidence and negative $p$ reflecting over-confidence. Note that one disadvantage of ACE is that it does not penalize excessively wide intervals (e.g., under-confident intervals that always cover $s_{i}$ incur a maximum $0.05$ penalty when $p=0.95$) as much as narrow ones. Hence, we treat ACE as a secondary, human-readable diagnostic and rely primarily on CRPS for PIE evaluation.
Figure 2: CE-Reg outperforms RF-LLMaaJ on RMSE, CRPS, Correlation, and ACE (95%) on Llama 3.2 11B (left) and Gemini 2.0 flash-lite (right), averaged across all datasets and task metrics. Lower is better for all except correlation. This result shows minimal difference between LLMs.
6 Findings
We now summarize 6 key empirical findings from evaluating PIE over 11 free-form generation tasks, 15 metrics, and 2 backbone LLMs. Rather than focus solely on point estimates or preference judgments, our experiments jointly assess the prediction of continuous evaluation scores and their associated uncertainty. We organize the results to study key properties of PIE, including feasibility, sample efficiency, metric-dependent behavior, and robustness across estimation techniques and models.
PIE is feasible Across datasets and target metrics, CE-based regression achieves an average CRPS of 0.10, RMSE of 0.16, and PC of 0.34 (Tables 3, 8, and 9), indicating that simple confidence features support useful prediction of continuous quality scores and calibrated uncertainty at inference time. This holds both for derived evaluation metrics and for human ratings (Table 4). Performance also stabilizes with as few as 16 labeled examples, suggesting PIE can be practical when annotation budgets are small.
@@ 6 Findings
PIE supports sample-efficient predictors As free-form gold references can be expensive to collect, it is important to understand how sample-efficient these methods would be. To answer that, we vary the size of the training set from as few as 4 instances and doubling up to the full training data size for CE-Reg. For RF-LLMaaJ, we limit up to 16 instances. CE-Reg achieves close to our best observed accuracy using few labeled examples: Figure 3 shows that CRPS converges after just 16 training instances across all datasets. Even when compared to RF-LLMaaJ’s maximum of 16 in-context examples (Table 5), CE-Reg delivers substantially lower error. Although performance varies slightly across datasets, the convergence point is consistent. This shows that the regression approach can be effectively trained with a small sample size, making it easy to collect training data even for out-of-domain tasks to mitigate potential domain differences.
PIE is more accurately addressed by CE-Reg than by RF-LLMaaJ. Because PIE requires predicting continuous evaluation metrics together with uncertainty at the instance level, not all evaluation paradigms are equally well matched to this objective. We therefore compare CE-Reg and RF-LLMaaJ as two representative paradigms for performance prediction under PIE.
Across datasets and evaluation metrics, CE-Reg consistently achieves substantially lower error under the primary CRPS objective than RF-LLMaaJ (Table 3). Averaged over task metrics, CE-Reg attains CRPS values averaging 0.11, compared to 0.14 for RF-LLMaaJ, indicating more informative predictive distributions. This advantage is also reflected in point accuracy: CE-Reg achieves RMSEs averaging 0.16, whereas RF-LLMaaJ incurs much larger errors, averaging around 0.22 (Table 8).
In addition, CE-Reg preserves stronger instance-level ordering, achieving consistently higher correlation of 0.28 across metrics than RF-LLMaaJ’s 0.09 (Table 9). Notably, for many task–metrics, RF-LLMaaJ exhibits near-zero correlation, indicating limited ability to recover instance-level signal. Increasing in-context examples from 4-16 yields negligible improvement (Table 5), suggesting that additional prompting alone does not reliably address this.
One plausible explanation is RF-LLMaaJ’s sensitivity to prompt formulation and target specification. Although both RF-LLMaaJ and verbalized confidence rely on in-context LLM predictions, verbalized confidence appears more stable across tasks, whereas RF-LLMaaJ must predict diverse task-specific metrics, which few-shot prompting may not reliably support. Taken together, these findings indicate that confidence estimates are informative signals for PIE but require supervision for effective calibration, while RF-LLMaaJ approaches remain less reliable across metrics and evaluation sources.
@@
PIE benefits most from graph-based consistency features. Among all CE-Reg, the four graph-based consistency measures (DegMat, Eccentricity, EigVal, LexSim) achieve the lowest CRPS, indicating that capturing variability across multiple generations provides a stronger predictive signal than one- or two-step verbalized confidence prompts, on all types of datasets and metrics. By contrast, the NumSemSets consistency metric performs on par with the verbalized methods, trailing the graph-based techniques. Better results from graph-based consistency over NumSemSets and verbalized confidence are more obviously reflected in the ACE and PC measures (Tables 10 and 9).
PIE yields consistent results across LLMs. Figure 2 evaluates the generality of CE-Reg and RF-LLMaaJ by comparing their performance across two backbone models. In addition to Llama 3.2 11B, we re-ran the same experimental setup on Gemini 2.0 Flash-Lite. Across both models and all evaluation criteria, we observe that CE-Reg outperforms RF-LLMaaJ, with absolute performance levels also comparable across models. This consistency suggests that the advantages of CE-Reg are not model-specific and that PIE generalizes across LLM families.
PIE performance varies across task metrics. Across datasets, BERTScore and CodeBLEU achieve low CRPS (Table 3) and high correlation (Table 9), indicating that their per-instance predictive distributions are accurate in absolute magnitude and that predicted scores are well-aligned in relative terms across instances. G-Eval metrics likewise attain low CRPS, but their correlation is consistently low, suggesting a different regime: predictions match typical score levels (often because true scores are tightly concentrated), yet are less informative for distinguishing instances.
In contrast, LLMEval, BiGGen Judge, GEMBA, and FactScore produce broader target distributions, leading to slightly higher CRPS while still maintaining relatively high correlation. This pattern suggests that these metrics contain more instance-to-instance variation in quality, enabling meaningful relative comparisons even when uncertainty-aware prediction is harder. Overall, different tasks and metrics exhibit different mixtures of (i) distributional predictability of absolute scores and (ii) relative, instance-discriminative predictability.
@@
Table 4: Direct Assessment: For BiGGen we evaluate two baselines using pre-generated outputs—Verbalized 2S and RF-LLMaaJ—for directly predicting human ratings. Methods requiring additional outputs are omitted. This result shows CE-Reg calibrating better than LLMaaJ on predicting human judgments.
@@
CRPS 0.09 0.09 0.09 0.14 0.14 0.14
Table 5: Comparing CE-Reg using DegMat feature vs. RF-LLMaaJ, varying training example count, averaged across all datasets, showing undetectable benefit from additional training examples on CRPS.
Figure 3: Impact of training data size on CE-Reg performance, measured by test-set CRPS across datasets using the selected task-specific metric (see Table 7). Training data sizes are capped at 64 instances. CRPS stabilizes after approximately 16 instances, suggesting a plateau for certain datasets.
@@ 7 Conclusion
In this work, we propose the task of Performance Interval Estimation (PIE) for predicting quality of free-form LLM generations. While confidence estimation is well-known for predicting the probability for whether model outputs are correct or not, predicting binary correctness in this way is most sensible when generation tasks have exact answers. In contrast, free-form generations are often more nuanced, with output quality being both fine-grained and multi-faceted.
We therefore motivate prediction of arbitrary continuous, reference-based metric scores (point estimates) with calibrated prediction intervals. Framing this prediction task as regression allows us to be agnostic as to both: 1) whether the prediction target is ordinal or continuous; and 2) whether we are predicting an observed value (e.g., human rating or program execution time) or a derived evaluation metric. We emphasize the importance of prediction intervals for risk-informed decision-making; a confidence interval lower-bounds expected quality to better assess the level of risk.
To promote this prediction task and foster community progress, we release our dataset, PIE-data, spanning 11 datasets across summarization, machine translation, code generation, function-calling, and question-answering, with multiple task evaluation metrics to predict. We also share our implemented methods and their outputs.
Our experiments compare two alternative prediction methods: confidence-based regression (CE-Reg), which extracts off-the-shelf verbalized and consistency-based confidence signals and trains classic regression models to output both point estimates and prediction intervals; and reference-free LLM-as-a-judge (RF-LLMaaJ), which prompts an LLM with in-context learning to yield metric estimates and 95% intervals for each input-output-instance. CE-Reg is substantially more accurate and better calibrated than RF-LLMaaJ, especially with consistency features, showing that lightweight supervised regression over simple confidence signals can outperform few-shot LLM judging for both prediction and uncertainty quantification.
@@ Acknowledgments
This research was supported in part by Cisco, Amazon, and by Good Systems [^3] https://goodsystems.utexas.edu/ (a UT Austin Grand Challenge dedicated to developing responsible AI technologies), as well as CosmicAI [^4] The NSF-Simons AI Institute for Cosmic Origins (CosmicAI): https://www.cosmicai.org/ support from the NSF (Cooperative Agreement 2421782) and the Simons Foundation (grant MPS-AI-00010515). We thank the Texas Advanced Computing Center (TACC) for use of its superb infrastructure. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
@@ A.2 Sample Prompts
      You will be given a news article. You will then be given one summary written for this article.      Your task is to rate the summary on one metric.      Please make sure you read and understand these instructions carefully.    Please keep this document open while reviewing, and refer to it as needed.      Evaluation Criteria:      {evaluation_instructions}      Evaluation Steps:      1. Read the news article carefully and identify the main facts and key points.    {internallinenumbers*} 2. Read the summary and compare it to the article. Check if the summary meets the {metric} criterion.    3. Assign a score for {metric} based on the Evaluation Criteria.      Source Text:      {source}      Summary:      {prediction}        Evaluation Form (scores ONLY), provide your response in the following format:    - {metric}:      evaluation_criteria = {    {internallinenumbers*} "Coherence": """Coherence (1-5) - the collective quality of all sentences. A well-structured summary should build from sentence to a coherent body of information about a topic.""",    "Consistency": """Consistency (1-5) - the factual alignment between the summary and the    {internallinenumbers*} source document. A factually consistent summary contains only statements entailed by the source document and avoids hallucinated facts.""",    {internallinenumbers*} "Fluency": """Fluency (1-3): the quality of the summary in terms of grammar, spelling, punctuation, word choice, and sentence structure.      - 1: Poor. The summary has many errors that make it hard to understand or sound unnatural.    {internallinenumbers*} - 2: Fair. The summary has some errors that affect the clarity or smoothness of the text, but the main points are still comprehensible.    - 3: Good. The summary has few or no errors and is easy to read and follow.""",    {internallinenumbers*} "Relevance": """Relevance (1-5) - selection of important content from the source. The summary should include only important information from the source document. Annotators were instructed to penalize summaries which contained redundancies and excess information."""    }
@@
      {internallinenumbers*} Task: Score the following LLM output of a question-answering task with respect to the following aspects using a 1 to 5 star rating system.      {internallinenumbers*} Dataset: The dataset is a Question-Answering dataset, specifically designed for evaluating factual precision and detailed comparative reasoning in AI-generated answers.      {internallinenumbers*} Output: Begin your evaluation by providing a short explanation. Be as objective as possible. After your explanation, provide your scores in JSON format like: [[SCORE]] {{"accuracy": 2, "informativeness": 3}}      Criteria:    - Accuracy:    1 star: Incorrect information    2 stars: Partially correct information    3 stars: Half correct information    4 stars: Mostly correct information    5 stars: Perfectly correct information      - Informativeness:    1 star: No or irrelevant information    2 stars: Very little information    3 stars: Some information    4 stars: Enough information    5 stars: Highly informative      Question:    {source}      Provided Answer:    {ans}      Reference Answer(s):    {ref_text}      Evaluation:    [[SCORE]] {{\"accuracy\": <1–5>, \"informativeness\": <1–5>}}
@@
      {internallinenumbers*} You are a fair judge assistant tasked with providing clear, objective feedback based on specific criteria, ensuring each assessment reflects the absolute standards set for performance.
@@
      ###Task Description:    {internallinenumbers*} An instruction (might include an Input inside it), a response to evaluate, and a score rubric representing an evaluation criteria are given.    {internallinenumbers*} 1. Write a detailed feedback that assesses the quality of the response strictly based on the given score rubric, not evaluating in general.    {internallinenumbers*} 2. After writing a feedback, write a score that is an integer between 1 and 5. You should refer to the score rubric.    {internallinenumbers*} 3. The output format should look as follows: "(write a feedback for criteria) [RESULT] (an integer number between 1 and 5)"    4. Please do not generate any other opening, closing, and explanations.      ###The instruction to evaluate:    {instruction}      ###Response to evaluate:    {response}      ###Score Rubrics:    {rubric}      ###Feedback:
@@
      {internallinenumbers*} You will be provided with a problem statement, a code snippet that supposedly addresses the problem,    and a catalog of code inconsistencies.    Evaluation Steps:    1. Read the problem statement carefully to identify the functionalities required for the    implementation.    {internallinenumbers*} 2. Read the code snippet and compare it to the problem statement. Check if the code snippet covers    the required functionalities.    3. Output your answer in a JSON format list.    {internallinenumbers*} a) If the code snippet is correct, output: ["inconsistency": "None", "severity": "Negligible"].    b) If the code snippet is incorrect, output the identified inconsistencies and their severity    {internallinenumbers*} according to the catalog of code inconsistencies. For example: ["inconsistency": "<inconsistency1>",    {internallinenumbers*} "severity": "<severity1>", "inconsistency": "<inconsistency2>", "severity": "<severity2>", ...]    Problem: {PROBLEM}    Code Snippet: {CODE}    Taxonomy of Common Inconsistencies:    1. Missing dependency declarations: Negligible    2. No error messages for unexpected input cases: Negligible    3. Inefficiency, unnecessary statements: Negligible    4. Edge case not handled: Small    5. Logic error: Major    6. Function or variable not defined: Fatal    7. Code not completed: Fatal    Evaluation Form:    JSON output (a JSON list only):